OpenStudy (anonymous):

simplifty completly 12x^3 - 4x^2 + 8x over -2x?

OpenStudy (callisto):

1. take out the common factor (12x^3 - 4x^2 + 8x ) / -2x = 4x (3x^2 - x +2) /-2x 2. cross out the common factor = -2(3x^2 - x +2) = -6x^2 +2x -4